Find the negative value for x such that 2^log base4(1/64) +x^2 =73/8
I would appreciate any help!
Thank you!

Answer by rothauserc(4718)

2^log base4(1/64) = 2^(-3.0) = 1/(2^3) = 1/8
1/8 +x^2 = 73/8
multiply 8 times both sides of =
1 +8x^2 = 73
8x^2 = 72
x^2 = 9
x = 3 or -3
you want the -3